# Spectral representations of interpolation spaces of reproducing kernel Hilbert spaces

### v1

@@ 3 Results
    The embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous with operator norm at most $\kappa$.
@@ 4 Application to statistical learning
Let us introduce some basic notions from statistical learning theory, where we refer to [16] for details. We assume that $(X,\mathcal{A})$ is a measurable space, and that $Y\subset\mathbb{R}$ is bounded. Suppose there is a probability measure $P$ on $X\times Y$, wherefrom only some samples $(x_{1},y_{1}),...,(x_{n},y_{n})$ are known. Typically, one searches for a prediction function $f:X\rightarrow\mathbb{R}$ that predicts an $Y$-value associated to a given $X$-value with low risk:
@@
where $P$ is in practice replaced by the empirical probability measure assigned to the given data. As done in [6], we assume in this section that the kernel of $H$ is bound by $1$.
Assuming that $L$ is the least squares loss, $f_{P,\lambda}$ can alternatively be found by the eigen-decomposition of $H$, see [18, Theorem 5.4]:
@@
where we assume that $P_{X}$ is the marginal distribution, that Y=[1,1]Y=[-1,1] and that $H$, the eigenvalues $\mu_{i}$ and the basis $e_{i}$ are as in Assumption H. By this reformulation, we can to describe various other methods of regularised minimisation than (19) by replacing the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ by a more general function $\mu_{i}g_{\lambda}(\mu_{i})$. The following conditions taken from [6] ensure that a reasonable algorithm emerges.
@@
Let $X$ be a measurable space, $P$ a measure on $X\times[-1,1]$ and $k$ a measurable kernel on $X$ with RKHS $H$ that is compactly embedded into $\mathrm{L}^{2}(P_{X})$, where the eigenvalues of $T_{k}$ are bounded by $1$. Let $0<\beta<1$, $0<\theta<1$ and $1\leq r\leq\infty$. If $f_{L,P}^{*}\in[\mathrm{L}^{2}(P_{X}),[H]_{\sim}]_{\beta,r}$, the following holds.
@@
Such estimates are play an important role in the statistical analysis of kernel-based learning algorithms, see e.g. [16, Theorem 7.23]. Note that the required existence of the embedding to $\mathrm{L}^{\infty}(\nu)$ may be checked via Theorem 3.3.
@@
    $=\lambda^{\gamma}\gamma^{\gamma}(1-\gamma)^{1-\gamma}$ (33)
@@
    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\lambda^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}$ (34)
@@
    $\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\gamma}\right)=\sup_{i\in I}(1-\mu_{i})^{t}\mu_{i}^{\gamma}\leq\sup_{0<\mu<1}(1-\mu)^{t}\mu^{\gamma}=\left(\frac{t}{\gamma+t}\right)^{t}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}$ (35)
@@ 5 Proofs
Let $(X,\mathcal{A},\nu)$ be a measure space, and $w:X\rightarrow[\epsilon,\infty)$ be measurable, where $\epsilon>0$ is some constant. Furthermore, let $s>1$, $\theta\in(0,1)$, and $r\in[1,\infty)$. Then
@@
The assertion directly follows from [7, Theorem 3.3], where we set $\mathscr{X}$ as $\mathrm{L}^{p}(\nu)$ and used the weight $\omega=w^{\frac{1}{p}}$, as the author applies the weight to the functions instead of the measures as we do here. Finally, we need to recall that the $K$-method and $J$-method used in [7] yield equivalent norms on the interpolation space. ∎
@@
$M_{j}:=\left\{x\in X\,\middle|\,\frac{1}{2^{j+1}}\leq\frac{1}{w(x)}<\frac{1}{2^{j}}\right\}\text{.}$ (38)
@@
Let $(X,\mathcal{A},\nu)$ be a measure space and let $\varepsilon>0$, and $w:X\rightarrow[\epsilon,\infty)$ be measurable. Furthermore, let $\theta\in(0,1)$, $r\in[1,\infty)$, and let
@@
Recall that $(\sqrt{\mu_{i}}e_{i})_{i}$ is an ONS in $H$. We write $H_{I}:=\overline{\operatorname{\mathrm{span}}\{\sqrt{\mu_{i}}e_{i}|i\in I\}}$ for the subspace spanned by this ONS. By [18, Formula (17)], we know that $H_{I}=\overline{\text{Im}I_{k}^{*}}$, and hence [18, Formula (19)] shows $H_{I}^{\perp}=\text{Ker}I_{k}$. If we choose an ONB $(\tilde{e}_{j})_{j}$ of $H_{I}^{\perp}$ on a new index set $J$, we thus find
$[H]_{\sim}$ $=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}+\sum_{j\in J}b_{i}\tilde{e}_{j}\right]_{\sim}\,\middle|\,(a_{i})\in\ell^{2}(I),(b_{j})\in\ell^{2}(J)\right\}$ (44)
$=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]\,\middle|\,(a_{i})\in\ell^{2}(I)\right\}$ (45)
@@
and thus $a_{i}\sqrt{\mu_{i}}=b_{i}$. Since $(a_{i})\in\ell^{2}(I)$, we get
@@
The map $\beta:\ell^{(\mu,\theta,r)}\rightarrow H_{\nu}^{(\theta,r)}$ given by $b\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ is an isometric isomorphism by Definition 2.2.
@@
By Lemma A.2, we know that $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$. Hence we can apply Theorem 3.1 to see that for the power spaces, we have $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$ as well. However, the operator norms of these two embedding operators might differ by a constant, as the norms are equivalent, but not necessarily equal.
@@
Let $0<\theta<1$ and $1\leq r\leq\infty$. If $b\in\ell^{(\mu,\theta,r)}$, then $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in the power space $H_{\nu}^{(\theta,r)}$.
@@
For finite $I$, the assertion is clear. Hence let us assume $I=\mathbb{N}$. Let $f:=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$, where convergence is in $\mathrm{L}^{2}(\nu)$ at first. To show convergence in $H_{\nu}^{(\theta,r)}$-norm, we notice that $\ell^{(\mu,\theta,r)}\triangleq[\ell^{2}(I),\allowbreak\ell^{2}(\mu^{-1})]_{(\theta,r)}$ by Lemma 5.2. Applying Lemma A.2 yields $\ell^{(\mu,\theta,r)}\hookrightarrow\ell^{2}(I)$. If $C$ denotes the corresponding operator norm, we thus find
$\left\|f-\sum_{i=1}^{n}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{(\mu,\theta,r)}\leq C\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{\ell^{2}(I)}\xrightarrow{n\rightarrow\infty}0\text{.}$ (61)
@@
which assigns to each $f$ the equivalence class of functions. This mapping is surjective by definition of $\mathrm{L}^{\infty}(\nu)$, hence it possesses a, not necessarily unique, right inverse. However, this reasoning does not yield further properties of such a right inverse. Nonetheless, by means of lifting theory, one can show that a bounded linear right inverse exists.
@@
The proof is based on lifting theory as developed in [8]. For a derivation of the specific result, see [18, Equation (58)].
So in other words, for each equivalence class in $\mathrm{L}^{\infty}(\nu)$, the operator $\phi$ chooses a bounded representative in a linear and continuous way.
Auxiliary results for the proofs of Theorems 3.2 and 3.3
Let us now introduce a setup which makes it possible to apply our results for direct sums of Banach spaces from Definition B.1.
@@
where we replaced $\pi^{(r)}(b)$ by $x$ and applied the isometry property on the right hand side. An application of Proposition B.5 shows that this is the case if and only if $(\pi_{j}(a))_{j\in\mathbb{Z}}\in\bigoplus_{j\in\mathbb{Z}}^{(r^{\prime})}E_{j}$ with norm $\|(\pi_{j}(a))_{j\in\mathbb{Z}}\|_{r^{\prime}}\leq C$. An application of the isometric isomorphism $\pi^{(r^{\prime})}$ finally shows that this is equivalent to $a\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|a\|_{(\mu,\theta,r^{\prime})}\leq C$, which is (ii). ∎
@@
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Let
$\overline{e}_{i}:=\phi([e_{i}]_{\sim})\text{,}$ (77)
where $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ is the lifting according to Theorem 5.7. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then the following statements are true.
@@
    It holds that $\sum_{i\in I}b_{i}\overline{e}_{i}=\sum_{i\in I}b_{i}e_{i}$ almost everywhere for all $b\in\ell^{(\mu,\theta,r)}$.
@@
    $H_{\nu}^{(\theta,r)}\rightarrow\overline{H}_{\nu}^{(\theta,r)},\quad\sum_{i\in I}b_{i}[e_{i}]_{\sim}\mapsto\sum_{i\in I}b_{i}\overline{e}_{i}$ (79)
@@
    $\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (80)

### v2

@@ 3 Results
    It holds that
    $\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa\text{.}$ (21)
  3. (iii)
    The embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous with operator norm at most $\kappa$.
Then we always have (i) $\Rightarrow$ (ii) $\Rightarrow$ (iii). Moreover, under Assumption H+, we have (i) $\Rightarrow$ (ii) $\Leftrightarrow$ (iii), and under Assumption T, we even have (i) $\Leftrightarrow$ (ii) $\Leftrightarrow$ (iii).
Remark 3.5.
Suppose that Assumption H is satisfied. Recall that the power kernel associated to $k$ is defined by its Mercer representation as
$k_{\nu}^{\theta}(x,x^{\prime})=\sum_{i\in I}\mu_{i}^{\theta}e_{i}(x)e_{i}(x^{\prime})$ (22)
for $x,x^{\prime}\in X$, see [27, Proposition 4.2]. Using this notion, we can reformulate the terms in (20) and (21) as follows.
$\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}=\left\|\sum_{i\in I}e_{i}(x)\mu_{i}^{\theta}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}$ $=\left\|\left[\sum_{i\in I}e_{i}(x)\mu_{i}^{\theta}e_{i}\right]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}$ (23)
$=\left\|\left[k_{\nu}^{\theta}(x,\cdot)\right]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}\text{.}$ (24)
Moreover, if $\sum_{i\in I}b_{i}e_{i}$ converges pointwise everywhere for all $b\in\ell^{(\mu,\theta,r^{\prime})}$, we may define $H_{\nu}^{(\theta,r^{\prime})}$ as in (17) and remove all $\cdot\,$$_{\sim}$ to consider functions instead of equivalence classes. Under Assumption T, this is possible due to Theorem 3.3.
Remark 3.6.
Suppose there is a $C>0$ such that the eigenfunctions $e_{i}$ satisfy $\|e_{i}\|_{\mathrm{L}^{\infty}(\nu)}<C$ for all $i\in I$. Then, by a countable union of null sets, we find that $|e_{i}(x)|<C$ for all $i\in I$ and for $\nu$-almost all $x\in X$. Then
$\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq C\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}\text{,}$ (25)
and hence condition (21) is satisfied for some $\kappa>0$ whenever $\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}<\infty$.
More explicitly, the term of interest $\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}$ is given by
$\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{r^{\prime}}{2}}\right)^{\frac{1}{r^{\prime}}}\quad\text{if }1\leq r^{\prime}<\infty\quad\text{ and by }\quad\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{1}{2}}\quad\text{if }r^{\prime}=\infty\text{.}$ (26)
Remark 3.7.
We illustrate the statements in Theorem 3.4 in a concrete situation. To this end, suppose that $\mu_{i}=i^{-\alpha}$ for $\alpha>1$, which is linked to Sobolev spaces [11]. Then we find for $j\in\mathbb{Z}$ that
$M_{j}=\{i\in\mathbb{N}\,|\,2^{j}<i^{-\alpha}\leq 2^{j+1}\}=\{i\in\mathbb{N}\,|\,2^{-\frac{j+1}{\alpha}}\leq i<2^{-\frac{j}{\alpha}}\}=[2^{-\frac{j+1}{\alpha}},2^{-\frac{j}{\alpha}})\cap\mathbb{N}\text{,}$ (27)
so $M_{j}=\{\}$ for $j\geq 0$. We hence note that $\min M_{j}\geq 2^{-\frac{j+1}{\alpha}}$ and $\max M_{j}\leq 2^{-\frac{j}{\alpha}}$. Furthermore, we have
$2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}-1\leq|M_{j}|\leq 2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}+1$ (28)
By that, we estimate for $j<0$
$\sum_{i\in M_{j}}i^{-1}\leq|M_{j}|2^{\frac{j+1}{\alpha}}\leq(2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}+1)2^{\frac{j+1}{\alpha}}\leq 2^{\frac{1}{\alpha}}-1+2^{\frac{j+1}{\alpha}}\leq 1+2^{\frac{j+1}{\alpha}}\leq 1+2^{\frac{1}{\alpha}}\text{.}$ (29)
Hence for $\theta=\frac{1}{\alpha}$ with $r=1$ and $r^{\prime}=\infty$, we get
$\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}=\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{1}{2}}=\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}i^{-1}\right)^{\frac{1}{2}}\leq\left(1+2^{\frac{1}{\alpha}}\right)^{\frac{1}{2}}<\infty\text{.}$ (30)
However, for $r=2$ with $r^{\prime}=2$, we get
$\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}=\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{r^{\prime}}{2}}\right)^{\frac{1}{r^{\prime}}}=\left(\sum_{j\in\mathbb{Z}}\sum_{i\in M_{j}}i^{-1}\right)^{\frac{1}{2}}=\left(\sum_{i\in\mathbb{N}}i^{-1}\right)^{\frac{1}{2}}=\infty\text{.}$ (31)
We see that in this situation, the fine index $r$ decides whether the conditions (20) and (21) are met.
@@ 4 Application to statistical learning
Let us introduce some basic notions from statistical learning theory, where we refer to [25] for details. We assume that $(X,\mathcal{A})$ is a measurable space, and that $Y\subset\mathbb{R}$ is bounded. Suppose there is a probability measure $P$ on $X\times Y$, wherefrom only some samples $(x_{1},y_{1}),...,(x_{n},y_{n})$ are known. Typically, one searches for a prediction function $f:X\rightarrow\mathbb{R}$ that predicts an $Y$-value associated to a given $X$-value with low risk
@@
where $P$ is in practice replaced by the empirical probability measure assigned to the given data. As done in [8], we assume in this section that the kernel of $H$ is bounded by $1$.
Assuming that $L$ is the least squares loss, $f_{P,\lambda}$ can alternatively be found by the eigen-decomposition of $H$, see [27, Theorem 5.4]:
@@
where we assume that $P_{X}$ is the marginal distribution, that Y=[1,1]Y=[-1,1] and that $H$, the eigenvalues $\mu_{i}$ and the basis $e_{i}$ are as in Assumption H. By this reformulation, we can describe various other methods of regularised minimisation than (33) by replacing the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ with a more general function $\mu_{i}g_{\lambda}(\mu_{i})$. The following conditions taken from [8] ensure that a reasonable algorithm emerges.
@@
Let $X$ be a measurable space, $P$ a measure on $X\times[-1,1]$ and $k$ a measurable kernel on $X$ with RKHS $H$ that is compactly embedded into $\mathrm{L}^{2}(P_{X})$, where the eigenvalues of $T_{k}$ are bounded by $1$. Let $0<\beta<1$, $0<\theta<1$ and $1\leq r\leq\infty$. If $f_{L,P}^{*}\in[\mathrm{L}^{2}(P_{X}),[H]_{\sim}]_{\beta,r}$, the following holds:
@@
Such estimates are play an important role in the statistical analysis of kernel-based learning algorithms, see e.g. [25, Theorem 7.23]. Note that the required existence of the embedding to $\mathrm{L}^{\infty}(\nu)$ may be checked via Theorem 3.4.
@@
    $=\lambda^{\gamma}\gamma^{\gamma}(1-\gamma)^{1-\gamma}\text{,}$ (47)
@@
    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\lambda^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{,}$ (48)
@@
    $\sup_{i\in I}\left(\big|1-\mu_{i}g_{\lambda}(\mu_{i})\big|\mu_{i}^{\gamma}\right)=\sup_{i\in I}(1-\mu_{i})^{t}\mu_{i}^{\gamma}\leq\sup_{0<\mu<1}(1-\mu)^{t}\mu^{\gamma}=\left(\frac{t}{\gamma+t}\right)^{t}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}\text{,}$ (49)
@@ 5 Proofs
Let $(X,\mathcal{A},\nu)$ be a measure space, and $w:X\rightarrow[\varepsilon,\infty)$ be measurable, where $\varepsilon>0$ is some constant. Furthermore, let $s>1$, $0<\theta<1$, and $1\leq r<\infty$. Then
@@
The assertion directly follows from [9, Theorem 3.3], where we set $\mathscr{X}$ as $\mathrm{L}^{p}(\nu)$ and use the weight $\omega=w^{\frac{1}{p}}$, as the author applies the weight to the functions instead of the measures as we do here. Finally, we need to recall that the $K$-method and $J$-method used in [9] yield equivalent norms on the interpolation space. ∎
@@
$M_{j}:=\left\{x\in X\,\middle|\,\frac{1}{2^{j+1}}\leq\frac{1}{w(x)}<\frac{1}{2^{j}}\right\}$ (52)
@@
Let $(X,\mathcal{A},\nu)$ be a measure space, let $\varepsilon>0$, and $w:X\rightarrow[\varepsilon,\infty)$ be measurable. Furthermore, let $0<\theta<1$, $1\leq r<\infty$, and let
@@
Recall that $(\sqrt{\mu_{i}}e_{i})_{i}$ is an ONS in $H$. We write $H_{I}:=\overline{\operatorname{\mathrm{span}}\{\sqrt{\mu_{i}}e_{i}|i\in I\}}$ for the subspace spanned by this ONS. By [27, Formula (17)], we know that $H_{I}=\overline{\text{Im}I_{k}^{*}}$, and hence [27, Formula (19)] shows $H_{I}^{\perp}=\text{Ker}I_{k}$. If we choose an ONB $(\tilde{e}_{j})_{j}$ of $H_{I}^{\perp}$ on a new index set $J$, we thus find
$[H]_{\sim}$ $=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}+\sum_{j\in J}b_{j}\tilde{e}_{j}\right]_{\sim}\,\middle|\,a\in\ell^{2}(I),b\in\ell^{2}(J)\right\}$ (58)
$=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]\,\middle|\,a\in\ell^{2}(I)\right\}$ (59)
@@
and thus $a_{i}\sqrt{\mu_{i}}=b_{i}$. Since $a\in\ell^{2}(I)$, we get
@@
The map $\beta:\ell^{(\mu,\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)}$ given by $b\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ is an isometric isomorphism by Definition 2.3.
@@
By Lemma A.2, we know that $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$. Hence we can apply Theorem 3.1 to see that for the power spaces, we have $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$ as well. However, the operator norms of these two embedding operators might differ by a constant, as the norms are equivalent, but not necessarily equal.
@@
Let $0<\theta<1$ and $1\leq r\leq\infty$. If $b\in\ell^{(\mu,\theta,r)}$, then $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in the power space $H]_{\sim}^{(\theta,r)}$.
@@
For finite $I$, the assertion is clear. Hence let us assume $I=\mathbb{N}$. Let $f:=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$, where convergence is in $\mathrm{L}^{2}(\nu)$ at first. To show convergence in $H]_{\sim}^{(\theta,r)}-$norm, we notice that $\ell^{(\mu,\theta,r)}\triangleq[\ell^{2}(I),\allowbreak\ell^{2}(\mu^{-1})]_{(\theta,r)}$ by Lemma 5.2. Applying Lemma A.2 yields $\ell^{(\mu,\theta,r)}\hookrightarrow\ell^{2}(I)$. If $C$ denotes the corresponding operator norm, we thus find
$\left\|f-\sum_{i=1}^{n}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{(\mu,\theta,r)}\leq C\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{\ell^{2}(I)}\xrightarrow{n\rightarrow\infty}0\text{.}$ (75)
@@
which assigns to each $f$ its equivalence class of functions. This mapping is surjective by definition of $\mathrm{L}^{\infty}(\nu)$, hence it possesses a, not necessarily unique, right inverse. However, this reasoning does not yield further properties of such a right inverse. Nonetheless, by means of lifting theory, one can show that a bounded linear right inverse exists.
@@
The proof is based on lifting theory as developed in [12]. For a derivation of the specific result, see [27, Equation (58)].
In other words, for each equivalence class in $\mathrm{L}^{\infty}(\nu)$, the operator $\phi$ chooses a bounded representative in a linear and continuous way.
Proofs of Theorems 3.2, 3.3, and 3.4
Let us now introduce a setup which makes it possible to apply the results for direct sums of Banach spaces from Appendix B.
@@
where we replaced $\pi^{(r)}(b)$ by $x$ and applied the isometry property on the right-hand side. An application of Proposition B.5 shows that this is the case if and only if $(\pi_{j}(a))_{j\in\mathbb{Z}}\in\bigoplus_{j\in\mathbb{Z}}^{(r^{\prime})}E_{j}$ with norm $\|(\pi_{j}(a))_{j\in\mathbb{Z}}\|_{r^{\prime}}\leq C$. An application of the isometric isomorphism $\pi^{(r^{\prime})}$ finally shows that this is equivalent to $a\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|a\|_{(\mu,\theta,r^{\prime})}\leq C$, which is (ii). ∎
In order to prove Theorem 3.2, we will use liftings from Theorem 5.7. As this requires a complete measure space, which is not necessarily given in Theorem 3.2, we consider the following lemma.
Lemma 5.11.
[…]
  1. (i)
    For every $\hat{\mathcal{A}}-$measurable $f:X\rightarrow\mathbb{R}$, there exists a $\mathcal{\nu}$-null set $N$ such that $\mathbbm{1}_{X\setminus N}f$ is $\mathcal{A}$-measurable.
  2. (ii)
    The operator
    $\mathrm{L}^{\infty}(\nu)\rightarrow\mathrm{L}^{\infty}(\hat{\nu}),\quad[f]_{\sim}\mapsto[f]_{\hat{\sim}}$ (91)
    is a well-defined isometric isomorphism, where $\cdot\,$$_{\hat{\sim}}$ denotes equivalence classes with respect to $\hat{\nu}$.
Proof.
(i). By [6, Proposition 2.2.5], there are $\mathcal{A}$-measurable functions $f_{0}$ and $f_{1}$ on $X$ with $f_{0}\leq f\leq f_{1}$ and $f_{0}=f_{1}$ $\nu$-almost everywhere. Choosing $N=\{x\in X\,|\,f_{0}(x)\neq f_{1}(x)\}$ shows that $\mathbbm{1}_{X\setminus N}f=\mathbbm{1}_{X\setminus N}f_{0}$ is $\mathcal{A}$-measurable.
(ii). We first prove that the mapping is well-defined. As $\mathcal{A}\subset\hat{\mathcal{A}}$, we find that any $\mathcal{A}$-measurable $f:X\rightarrow\mathbb{R}$ is $\hat{\mathcal{A}}-$measurable, which justifies writing $f]_{\hat{\sim}}$. To see that the choice of representative of $f]_{\sim}$ is irrelevant, let $f$ and $g$ be $\mathcal{A}$-measurable functions with $f]_{\sim}=[g]_{\sim}$. Then $f-g]_{\sim}=0$, hence there is a $\nu$-null set $N$ such that f(x)g(x)=0f(x)-g(x)=0 for all $x\in X\setminus N$. But by definition of the completion, $N$ is a $\hat{\nu}-$zero set as well, which shows that $f-g]_{\hat{\sim}}=0$. Hence we get $f]_{\hat{\sim}}=[g]_{\hat{\sim}}$.
The linearity of the mapping is obvious. Moreover, for an $\mathcal{A}$-measurable $f:X\rightarrow\mathbb{R}$ with $f]_{\sim}\in\mathrm{L}^{\infty}(\nu)$, we have
$\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$ $=\inf\{B\geq 0\,|\,\nu(\{x\in X\,|\,f(x)>B\})=0\}$ (92)
$=\inf\{B\geq 0\,|\,\hat{\nu}(\{x\in X\,|\,f(x)>B\})=0\}$ (93)
$=\|[f]_{\hat{\sim}}\|_{\mathrm{L}^{\infty}(\hat{\nu})}\text{,}$ (94)
[…]
@@
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then for all $i\in I$, there exist representatives $\overline{e}_{i}\in\mathcal{L}^{\infty}(X)$ of $e_{i}]_{\sim}$ such that the following statements are true:
@@
    There is a $\nu$-null set $N\subset X$ such that $\sum_{i\in I}b_{i}\overline{e}_{i}(x)=\sum_{i\in I}b_{i}e_{i}(x)$ for all $b\in\ell^{(\mu,\theta,r)}$ and all $x\in X\setminus N$.
@@
    $[H]_{\sim}^{(\theta,r)}\rightarrow\overline{H}_{\nu}^{(\theta,r)},\quad\sum_{i\in I}b_{i}[e_{i}]_{\sim}\mapsto\sum_{i\in I}b_{i}\overline{e}_{i}$ (96)
@@
    $\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (97)